# UGG-ReID: Uncertainty-Guided Graph Model for Multi-Modal Object Re-Identification

### v2

@@ Abstract
Multi-modal object Re-IDentification (ReID) has gained considerable attention with the goal of retrieving specific targets across cameras using heterogeneous visual data sources. Existing methods primarily aim to improve identification performance, but often overlook the uncertainty arising from inherent defects, such as intra-modal noise and inter-modal conflicts. This uncertainty is particularly significant in the case of fine-grained local occlusion and frame loss, which becomes a challenge in multi-modal learning. To address the above challenge, we propose a robust approach named Uncertainty-Guided Graph model for multi-modal object ReID (UGG-ReID). UGG-ReID is designed to mitigate noise interference and facilitate effective multi-modal fusion by estimating both local and sample-level aleatoric uncertainty and explicitly modeling their dependencies. Specifically, we first propose the Gaussian patch-graph representation model that leverages uncertainty to quantify fine-grained local cues and capture their structural relationships. This process boosts the expressiveness of modal-specific information, ensuring that the generated embeddings are both more informative and robust. Subsequently, we design an uncertainty-guided mixture of experts strategy that dynamically routes samples to experts exhibiting low uncertainty. This strategy effectively suppresses noise-induced instability, leading to enhanced robustness. Meanwhile, we design an uncertainty-guided routing to strengthen the multi-modal interaction, improving the performance. UGG-ReID is comprehensively evaluated on five representative multi-modal object ReID datasets, encompassing diverse spectral modalities. Experimental results show that the proposed method achieves excellent performance on all datasets and is significantly better than current methods in terms of noise immunity. Our code will be made public upon acceptance.
@@ 1 Introduction
Multi-modal data has emerged as a significant trend in artificial intelligence [1, 2, 3, 4]. Especially driven by large model technology, more and more applications have begun to utilize multi-modal information for comprehensive analysis [5, 6, 7]. Multi-modal object Re-IDentification (ReID) [8, 9, 10, 11], as a cutting-edge direction of this research, not only broadens the application boundaries of traditional object ReID [12, 13], but also effectively makes up for some limitations in cross-modal object ReID [14, 15, 16].
@@
Recently, researchers have extensively explored multi-modal feature fusion and matching strategies to bridge the representation gap between different modalities for object ReID [9, 17, 18, 19]. For example, Zheng et al. [9] propose a progressive fusion method to achieve effective fusion of multi-modal data. Wang et al. [17] propose HTT that exploits the relationship on unseen test data between heterogeneous modalities to improve performance. On the other hand, existing methods [20, 21, 22, 23] have begun to pay attention to the effect of modal noise on the discriminative properties of local regions, such as the introduction of local region alignment, noise suppression module or sample reconstruction, to enhance the robustness and improve the performance of the model. Among these, Zhang et al. [21] propose EDITOR to suppress interference from background information and promote feature learn. Wang et al. [22] propose utilizing CDA to focus on localized regions of the discriminatory. We can observe that some methods usually assume that the quality and representation of each modal data are balanced and stable, and ignore the local and sample noise disturbances within the modality caused by factors such as occlusion, low resolution [9, 17]. Although existing works [21, 22] have achieved some success in mitigating the impact of noise on network performance, there are still significant shortcomings in their approaches when confronted with inconsistent local noise patterns and samples with different noise intensities. As depicted in Fig. 1 (a), the lack of fine-grained local noise learn and sample-level noise handling compromises the robustness of multi-modal fusion, which increases multi-modal uncertainty and impairs overall model performance.
@@
Therefore, to solve the above problems, we provide a robust approach named Uncertainty-Guided Graph model for multi-modal object ReID (UGG-ReID). The proposed method quantifies local aleatoric uncertainty and models their structural dependencies. Following this, leveraging per-sample uncertainty guides feature refinement, promoting more reliable modality interaction. To be specific, we first propose the Gaussian Patch-Graph Representation (GPGR), which encourages fine-grained local features to conform to Gaussian distributions. Meanwhile, a Gaussian patch graph is constructed to explicitly model the dependencies among these local features, thereby capturing fine-grained consistency information. Note that previous methods typically adopt fixed strategies to extract fine-grained local features [21, 22]. In contrast, benefiting from the idea of uncertainty modeling [24, 25, 26], we represent local regions as Gaussian distributions to explicitly quantify uncertainty and employ relational modeling to capture dependencies between local cues. This approach facilitates more effective learning of the diversity inherent in fine-grained local information, resulting in richer representations and increased robustness. As illustrated in Fig. 1 (b), the uncertainties of local tokens are effectively evaluated and explored via uncertainty-based modeling.
Second, we design an Uncertainty-Guided Mixture of Experts (UGMoE) strategy, which dynamically assigns samples to different experts based on their uncertainty levels. Besides, this strategy incorporates an uncertainty-guided routing mechanism to enhance multi-modal interaction. Unlike the existing methods, which only utilize Mixture-of-Experts (MoE) to promote multi-modal fusion, the differences between different modalities of the sample are ignored [27, 28]. As shown in Fig. 1 (c), multi-modal data is jointly learned to improve the overall performance based on uncertainty. The entire framework is trained in an end-to-end manner. Through extensive experimentation on five public multi-modal object ReID datasets, our method not only achieves competitive performance but also consistently outperforms prior methods in noisy scenarios, validating its effectiveness and robustness.
@@ 2.1 Multi-modal Object ReID
The existing multi-modal ReID methods can be summarized into two types: one focuses on feature fusion and integration between modalities, aiming to alleviate the semantic representation differences between different modalities [18, 23, 29, 30]. For instance, Yang et al. [30] propose the tri-interaction enhancement network (TIENet). This method applies spatial-frequency interaction to enhance feature extraction and multi-modal fusion. Wang et al. [18] propose a novel method called MambaPro which utilizes mamba aggregation to complementary information of multi-modal Object ReID. Zhang et al. [23] propose PromptMA to establish effective connections among different multi-modal information. The other type focuses on the local noise interference within the modality and improves the overall recognition performance by enhancing the local region discrimination [10, 20, 21, 22]. Zheng et al. [10] proposes CCNet, which seeks to simultaneously mitigate identification uncertainty due to modal variability and sample appearance changes by jointly modeling multi-modal heterogeneity and intraclass perturbations under view angle and lighting changes. Zhang et al. [21] propose EDITOR that uses Spatial-Frequency Token Selection (SFTS) module to select diverse tokens and suppress the effects of background interference. Wang et al. [22] propose the Inverted text with cooperative DEformable Aggregation (IDEA) framework to solve noise interference, enhancing feature robustness. However, the existing methods generally lack explicit modeling of fine-grained local information quality and sample-level uncertainty, which makes it difficult for the model to effectively perceive and suppress local noise in the face of complex modal degradation or multi-source noise interference, which limits its robustness and generalization ability.
@@ 2.2 Uncertainty in Multi-Modal Learning
Multi-modal learning faces the challenge of uncertainty from the data layer and the model layer [1, 31, 32, 24, 25, 26]. The former is embodied in aleatoric uncertainty, which arises from indelible perceived noise; The latter manifests as an epistemic uncertainty due to the limitations of the model’s capabilities. In practice, Aleatoric uncertainty is more common and has a direct impact on model performance. To this end, many works in recent years have focused on the introduction of uncertainty mechanisms to multi-modal learning to effectively identify and suppress unreliable information, improving the performance of the model. Ji et al. [31] propose the Probability Distribution Encoder (PDA) module to aggregate all modalities into a probability distribution, framing the uncertainty and optimizing multi-modal representations. Gao et al. [1] propose quantifying the intrinsic aleatoric uncertainty of single modality to enhance multi-modal features. Zhang et al. [26] propose UMLMC that employs uncertainty-guided meta-learning to mitigate feature-level bias. Therefore, reasonable modeling and quantification of the uncertainty in multi-modal data can not only improve the robustness and generalization ability of multi-modal models but also provide more reliable decision support for practical applications.
@@ 3 Methodology
We propose a novel Uncertainty-Guided multi-modal object ReID (UGG-ReID) framework, as shown in Fig. 2. Specifically, the proposed UGG-ReID first designs the Gaussian Patch-Graph Representation (GPGR) model as Fig. 2 (a), which quantifies both global and local uncertainties and models the dependencies between them. This enables richer feature representations for each modality and yields semantically more consistent embeddings. Subsequently, the Uncertainty-Guided Mixture of Experts (UGMoE) strategy makes samples select experts of low uncertainty, promoting multi-modal interactions while mitigating the propagation of excessive noise. Overall, this framework injects a controlled amount of sample noise during the learning process and effectively captures multi-modal information, thereby enhancing the model’s robustness and performance. Below, we will introduce each component of UGG-ReID in detail.
@@ 3.1 Feature Initialization
To align with prior research [27, 18], we adopt the visual encoder of CLIP with a shared backbone to extract initial features $x^{m}=\{x^{m}_{c},x^{m}_{p}\}$ of multi-modal data. Here, $m\in\{R,N,T\}$ indexes the RGB, Near-Infrared (NIR), and Thermal Infrared (TIR) modalities, while $x^{m}_{c}$ and $x^{m}_{p}=\{x^{m}_{1},x^{m}_{2}\cdots x^{m}_{n}\}$ denote class tokens and local tokens, respectively. $n$ is the number of local tokens.
@@ 3.2.1 Gaussian Patch-Graph Learning
where $\text{FC}_{\mu}$ and $\text{FC}_{\sigma}$ are projection layers to learn the mean and standard deviation, respectively. We denote the means and variances by ${\boldsymbol{\mu}^{m}}=\{{\mu}^{m}_{c},{\mu}^{m}_{1},{\mu}^{m}_{2}\cdots{\mu}^{m}_{n}\}\in\mathbb{R}^{N\times D}$ and $\boldsymbol{\sigma}^{m}=\{{\sigma}_{c}^{m},{\sigma}_{1}^{m},{\sigma}_{2}^{m}\cdots{\sigma}_{n}^{m}\}\in\mathbb{R}^{N\times D}$, respectively. Then, we calculate the edge weight in $E^{m}$ and construct the structural relationship to achieve global and local information interaction. Let $\boldsymbol{\alpha}^{m}$ be the adjacency matrix and encode structural relationships of $m$-th modality. Specifically, we compute the similarity by utilizing the mean vectors ${\boldsymbol{\mu}^{m}}$ to avoid the interference of highly uncertain nodes,
@@
where ${\mu}_{i}^{m}$ and ${\mu}_{j}^{m}$ are the mean representations of $i$-th and $j$-th nodes in $m$-th modality. $\mathrm{Similarity}$ is a metric function and we adopt Euclidean distance [33] in our experiments.
@@ 3.2.2 Gaussian Patch-Graph Convolutional Network
After constructing the Gaussian Patch-Graph, we further employ a Gaussian Patch-Graph Convolutional Network (GPGCN) to facilitate message passing among distributional nodes. Since each node is modeled as a Gaussian distribution, traditional graph convolution operations are not directly applicable. Thus, we adopt GPGCN to separately conduct message passing based on the mean and variance and thus enable effective propagation of both semantic representations and uncertainty information. Formally, the message passing process is defined as,
@@
where $\boldsymbol{\hat{\mu}}^{(m,0)}=\boldsymbol{\mu}^{m}$ and $\boldsymbol{\hat{\sigma}}^{(m,0)}=\boldsymbol{\sigma}^{(m)}$. $l=0,1\cdots L-1$ denotes the $l$-th layers of GGCN and $\mathcal{D}^{m}$ is corresponding degree matrix of $\boldsymbol{\alpha}^{m}$. $\boldsymbol{w_{\hat{\mu}}}^{(m,l)}$ and $\boldsymbol{w_{\hat{\sigma}}}^{(m,l)}$ are trainable weight matrices. After stacking multiple layers, we can obtain the final output $\boldsymbol{\hat{\mu}}^{(m,L)}$ and $\boldsymbol{\hat{\sigma}}^{(m,L)}$. For convenience, let $\boldsymbol{\hat{\mu}}^{m}=\boldsymbol{\hat{\mu}}^{(m,L)}$ and $\boldsymbol{\hat{\sigma}}^{m}=\boldsymbol{\hat{\sigma}}^{(m,L)}$ in the following text. Considering that node representations follow Gaussian distributions, we adopt the sampling operation to obtain the final output, i.e., $\boldsymbol{\hat{x}}^{m}\sim\mathcal{N}(\boldsymbol{\hat{\mu}}^{m},\boldsymbol{\hat{\sigma}^{2}}^{m})$. However, this sampling process is not differentiable. Therefore, we employ the reparameterization trick to generate node representations as follows,
$\boldsymbol{\hat{x}}^{m}=\boldsymbol{\hat{\mu}}^{m}+\epsilon\boldsymbol{\hat{\sigma}}^{m},\quad\epsilon\sim\mathcal{N}(0,I),$ (4)
@@
where $\eta$ denotes a pooling operation. [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] is the concatenation operation and $W^{m}$ is the transformation matrix. To further promote joint aggregation representation, Kullback-Leibler divergence [34] is applied to the class token as,

### v3

@@ Abstract
Multi-modal object Re-IDentification (ReID) has gained considerable attention with the goal of retrieving specific targets across cameras using heterogeneous visual data sources. At present, multi-modal object ReID faces two core challenges: (1) learning robust features under fine-grained local noise caused by occlusion, frame loss, and other disruptions; and (2) effectively integrating heterogeneous modalities to enhance multi-modal representation. To address the above challenges, we propose a robust approach named Uncertainty-Guided Graph model for multi-modal object ReID (UGG-ReID). UGG-ReID is designed to mitigate noise interference and facilitate effective multi-modal fusion by estimating both local and sample-level aleatoric uncertainty and explicitly modeling their dependencies. Specifically, we first propose the Gaussian patch-graph representation model that leverages uncertainty to quantify fine-grained local cues and capture their structural relationships. This process boosts the expressiveness of modal-specific information, ensuring that the generated embeddings are both more informative and robust. Subsequently, we design an uncertainty-guided mixture of experts strategy that dynamically routes samples to experts exhibiting low uncertainty. This strategy effectively suppresses noise-induced instability, leading to enhanced robustness. Meanwhile, we design an uncertainty-guided routing to strengthen the multi-modal interaction, improving the performance. UGG-ReID is comprehensively evaluated on five representative multi-modal object ReID datasets, encompassing diverse spectral modalities. Experimental results show that the proposed method achieves excellent performance on all datasets and is significantly better than current methods in terms of noise immunity. Our code is available at https://github.com/wanxixi11/UGG-ReID.
@@ 1 Introduction
Multi-modal data has emerged as a significant trend in artificial intelligence [10657706, 10655615, 10.1145/3696410.3714606, 10841969]. Especially driven by large model technology, more and more applications have begun to utilize multi-modal information for comprehensive analysis [mckinzie2024mm1, 10655344, 10.5555/3692070.3694257]. Multi-modal object Re-IDentification (ReID) [Li2020MultiSpectralVR, Zheng2021RobustMP, ZHENG2023101901, ZHENG2025102800], as a cutting-edge direction of this research, not only broadens the application boundaries of traditional object ReID [9011001, 10807364, s44267-023-00032-9], but also effectively makes up for some limitations in cross-modal object ReID [Dai2018CrossModalityPR, 10.1145/3343031.3351006, He_2021_ICCV].
@@
Recently, researchers have extensively explored multi-modal feature fusion and matching strategies to bridge the representation gap between different modalities for object ReID [Zheng2021RobustMP, Wang2024HeterogeneousTT, Wang2024MambaPro, Yang2025TIENet]. For example, Zheng et al. [Zheng2021RobustMP] propose a progressive fusion method to achieve effective fusion of multi-modal data. Wang et al. [Wang2024HeterogeneousTT] propose HTT that exploits the relationship on unseen test data between heterogeneous modalities to improve performance. On the other hand, existing methods [Wang2023TOPReIDMO, 10654953, wang2025IDEA, Zhang25PromptMA] have begun to pay attention to the effect of modal noise on the discriminative properties of local regions, such as the introduction of local region alignment, noise suppression module or sample reconstruction, to enhance the robustness and improve the performance of the model. Among these, Zhang et al. [10654953] propose EDITOR to suppress interference from background information and promote feature learn. Wang et al. [wang2025IDEA] propose utilizing CDA to focus on localized regions of the discriminatory. We can observe that some methods usually assume that the quality and representation of each modal data are balanced and stable, and ignore the local and sample noise disturbances within the modality caused by factors such as occlusion, low resolution [Zheng2021RobustMP, Wang2024HeterogeneousTT]. Although existing works [10654953, wang2025IDEA, YangTwinNoisy, YangRobust, Qin2023NoisyCorrespondenceLF] have achieved some success in mitigating the impact of noise on network performance, there are still significant shortcomings in their approaches when confronted with inconsistent local noise patterns and samples with different noise intensities. As depicted in Fig. 1 (a), the lack of fine-grained local noise learn and sample-level noise handling compromises the robustness of multi-modal fusion, which increases multi-modal uncertainty and impairs overall model performance.
@@
Therefore, to solve the above problems, we provide a robust approach named Uncertainty-Guided Graph model for multi-modal object ReID (UGG-ReID). The proposed method quantifies local aleatoric uncertainty and models their structural dependencies. Following this, leveraging per-sample uncertainty guides feature refinement, promoting more reliable modality interaction. To be specific, we first propose the Gaussian Patch-Graph Representation (GPGR), which encourages fine-grained local features to conform to Gaussian distributions. Meanwhile, a Gaussian patch graph is constructed to explicitly model the dependencies among these local features, thereby capturing fine-grained consistency information. As illustrated in Fig. 1 (b), the uncertainties of local tokens are effectively evaluated and explored via uncertainty-based modeling. Second, we design an Uncertainty-Guided Mixture of Experts (UGMoE) strategy, which dynamically assigns samples to different experts based on their uncertainty levels. Besides, this strategy incorporates an uncertainty-guided routing mechanism to enhance multi-modal interaction. As shown in Fig. 1 (c), multi-modal data is jointly learned to improve the overall performance based on uncertainty. The entire framework is trained in an end-to-end manner. Through extensive experimentation on five public multi-modal object ReID datasets, our method not only achieves competitive performance but also consistently outperforms prior methods in noisy scenarios, validating its effectiveness and robustness.
@@ 2.1 Multi-Modal Object ReID
The existing multi-modal ReID methods can be summarized into two types: one focuses on feature fusion and integration between modalities, aiming to alleviate the semantic representation differences between different modalities [Wang2024MambaPro, Zhang25PromptMA, li2025icpl, Yang25TIENet]. For instance, Yang et al. [Yang25TIENet] propose the tri-interaction enhancement network (TIENet). This method applies spatial-frequency interaction to enhance feature extraction and multi-modal fusion. Wang et al. [Wang2024MambaPro] propose a novel method called MambaPro, which utilizes mamba aggregation to fuse the information of multi-modal Object ReID. Zhang et al. [Zhang25PromptMA] propose PromptMA to establish effective connections among different multi-modal information. The other type focuses on the local noise interference within the modality and improves the overall recognition performance by enhancing the local region discrimination [ZHENG2023101901, Wang2023TOPReIDMO, 10654953, wang2025IDEA]. Zheng et al. [ZHENG2023101901] proposes CCNet, which seeks to simultaneously mitigate identification uncertainty due to modal variability and sample appearance changes by jointly modeling multi-modal heterogeneity and intraclass perturbations under view angle and lighting changes. Zhang et al. [10654953] propose EDITOR that uses Spatial-Frequency Token Selection (SFTS) module to select diverse tokens and suppress the effects of background interference. Wang et al. [wang2025IDEA] propose the Inverted text with cooperative DEformable Aggregation (IDEA) framework to solve noise interference, enhancing feature robustness. However, the existing methods generally lack explicit modeling of fine-grained local information quality and sample-level uncertainty, which makes it difficult for the model to effectively perceive and suppress local noise in the face of complex modal degradation or multi-source noise interference, which limits its robustness and generalization ability.
@@ 2.2 Uncertainty in Multi-Modal Learning
Multi-modal learning faces the challenge of uncertainty from the data layer and the model layer [10657706, 10205473, 10.1145/3664647.3680949, 10737387, 10688376, ZHANG2025110993, s44267-025-00078-x]. The former is embodied in aleatoric uncertainty, which arises from indelible perceived noise; The latter manifests as an epistemic uncertainty due to the limitations of the model’s capabilities. In practice, aleatoric uncertainty is more common and has a direct impact on model performance. To this end, many works in recent years have focused on the introduction of uncertainty mechanisms to multi-modal learning to effectively identify and suppress unreliable information, improving the performance of the model. Ji et al. [10205473] propose the Probability Distribution Encoder (PDE) module to aggregate all modalities into a probability distribution, framing the uncertainty and optimizing multi-modal representations. Gao et al. [10657706] propose quantifying the intrinsic aleatoric uncertainty of single modality to enhance multi-modal features. Zhang et al. [ZHANG2025110993] propose UMLMC that employs uncertainty-guided meta-learning to mitigate feature-level bias. Therefore, reasonable modeling and quantification of the uncertainty in multi-modal data can not only improve the robustness and generalization ability of multi-modal models but also provide more reliable decision support for practical applications.
@@ 3 Methodology
We propose a novel Uncertainty-Guided multi-modal object ReID (UGG-ReID) framework, as shown in Fig. 2. Specifically, the proposed UGG-ReID first designs the Gaussian Patch-Graph Representation (GPGR) model as Fig. 2 (a), which quantifies both global and local uncertainties and models the dependencies between them. This enables richer feature representations for each modality and yields semantically more consistent embeddings. Subsequently, the Uncertainty-Guided Mixture of Experts (UGMoE) strategy makes samples select experts of low uncertainty, promoting multi-modal interactions while mitigating the propagation of excessive noise. Overall, this framework injects a controlled amount of sample noise during the learning process and effectively captures multi-modal information, thereby enhancing the model’s robustness and performance. Below, we will introduce each component of UGG-ReID with details.
@@ 3.1 Feature Initialization
To align with prior research [wang2024demo, Wang2024MambaPro], we adopt the visual encoder of CLIP with a shared backbone to extract initial features $x^{m}=\{x^{m}_{c},x^{m}_{p}\}$ of multi-modal data. Here, $m\in\{R,N,T\}$ indexes the RGB, Near‑Infrared (NIR), and Thermal Infrared (TIR) modalities, while $x^{m}_{c}$ and $x^{m}_{p}=\{x^{m}_{1},x^{m}_{2}\cdots x^{m}_{n}\}$ denote class tokens and local tokens, respectively. $n$ is the number of local tokens.
@@ 3.2.1 Gaussian Patch-Graph Learning
where $\text{FC}_{\mu}$ and $\text{FC}_{\sigma}$ are projection layers to learn the mean and standard deviation, respectively. We denote the means and variances by ${\boldsymbol{\mu}^{m}}=\{{\mu}^{m}_{c},{\mu}^{m}_{1},{\mu}^{m}_{2}\cdots{\mu}^{m}_{n}\}\in\mathbb{R}^{N\times D}$ and $\boldsymbol{\sigma}^{m}=\{{\sigma}_{c}^{m},{\sigma}_{1}^{m},{\sigma}_{2}^{m}\cdots{\sigma}_{n}^{m}\}\in\mathbb{R}^{N\times D}$, respectively. Then, we calculate the edge weight in $E^{m}$ and construct the structural relationship to achieve global and local information interaction. Let $\boldsymbol{\alpha}^{m}$ be the adjacency matrix and encode structural relationships of $m$-th modality. Specifically, we compute the similarity by utilizing the mean vectors $\boldsymbol{\mu}^{m}$} to avoid the interference of high uncertain nodes,
@@
where ${\mu}_{i}^{m}$ and ${\mu}_{j}^{m}$ are the mean representations of $i$-th and $j$-th nodes in $m$-th modality. $\mathrm{Similarity}$ is a metric function and we adopt Euclidean distance [10.1145/3459637.3482261] in our experiments.
@@ 3.2.2 Gaussian Patch-Graph Convolutional Network
After contructing the Gaussian Patch-Graph, we further employ a Gaussian Patch-Graph Convolutional Network (GPGCN) to facilitate message passing among distributional nodes. Since each node is modeled as a Gaussian distribution, traditional graph convolution operations are not directly applicable. Thus, we adopt GPGCN to separately conduct message passing based on the mean and variance [10.1145/3292500.3330851] and thus enable effective propagation of both semantic representations and uncertainty information. Formally, the message passing process is defined as,
@@
where $\boldsymbol{\hat{\mu}}^{(m,0)}=\boldsymbol{\mu}^{m}$ and $\boldsymbol{\hat{\sigma}}^{(m,0)}=\boldsymbol{\sigma}^{(m)}$. $l=0,1\cdots L-1$ denotes the $l$-th layers of GGCN and $\mathcal{D}^{m}$ is corresponding degree matrix of $\boldsymbol{\alpha}^{m}$. $\boldsymbol{w_{\hat{\mu}}}^{(m,l)}$ and $\boldsymbol{w_{\hat{\sigma}}}^{(m,l)}$ are trainable weight matrices. After stacking multiple layers, we can obtain the final output $\boldsymbol{\hat{\mu}}^{(m,L)}$ and $\boldsymbol{\hat{\sigma}}^{(m,L)}$. For convenience, let $\boldsymbol{\hat{\mu}}^{m}=\boldsymbol{\hat{\mu}}^{(m,L)}$ and $\boldsymbol{\hat{\sigma}}^{m}=\boldsymbol{\hat{\sigma}}^{(m,L)}$ in the following text. Considering that node representations follow Gaussian distributions, we adopt the sampling operation to obtain final output, i.e., $\boldsymbol{\hat{x}}^{m}\sim\mathcal{N}(\boldsymbol{\hat{\mu}}^{m},\boldsymbol{\hat{\sigma}^{2}}^{m})$. However, this sampling process is not differentiable. Therefore, we employ the reparameterization trick to generate node representations as follows,
$\boldsymbol{\hat{x}}^{m}=\boldsymbol{\hat{\mu}}^{m}+\epsilon\boldsymbol{\hat{\sigma}}^{m},\quad\epsilon\sim\mathcal{N}(0,I)$ (4)
@@
where $\eta$ denotes a pooling operation. $\cdot$$\cdot$] is the concatenation operation and $W^{m}$ is the transformation matrix. To further promote joint aggregation representation, Kullback-Leibler divergence [9157756] is applied to the class token as,